OpenStudy (anonymous):

On a journey of 600 km, a train was delayed 1h and 30 min after having covered 1/4 of the way. To arrive on time at the destination, the engine driver had to increase the speed by 15km/hr. How long did the train travel for?

OpenStudy (anonymous):

does that mean it reached 1/4 of its way in 1 hr 30 mins ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

it would a bit less than 5 hr and 30 mins.

ganeshie8 (ganeshie8):

lets say the train was travelling at \(v1\) kmph in the first quarter of journey

ganeshie8 (ganeshie8):

then, after 1/4th journey, it travels at \(v1 + 15\) kmph

ganeshie8 (ganeshie8):

we can have these equations :- \(\large v1(\frac{t}{4} + 1.5) = 150 ----------- (1)\) \(\large (v1+15)(\frac{3t}{4} - 1.5) = 450 ----------- (2)\)

ganeshie8 (ganeshie8):

two equations, two unknowns

ganeshie8 (ganeshie8):

you can solve them ?

OpenStudy (raden):

600/v = 150/v + 3/2 + 450/(v + 15) simplify becomes : (300 - v)/(2v) = 150/(v + 15) 300v = 4500 + 285v - v^2 v^2 - 15v - 4500 = 0 (v - 60)(v + 75) = 0 only v = 60 km/h which satifies therefore : t = d/v = 600/60 = 10 hours

OpenStudy (dan815):

haha vmath why do u always ask these kinds of algebra questions

OpenStudy (dan815):

ive seen u ask these kinds for months

OpenStudy (whpalmer4):

The slip in ganeshie8's approach is that the first equation gives "distance credit" for that the hour and a half delay spent moving at \(v_1\), but of course that isn't true, as the train isn't moving during the delay. It should be \[v_1(\frac{t}{4}) + 0(1.5) = 150\]and combined with the second equation will give the correct answer.